OpenStudy (anonymous):

slove plz help csc x- 2=0

OpenStudy (anonymous):

it needs more info...can u post the options if its a multipul choice question pls

OpenStudy (anonymous):

you just add 2 to both sides x-2=0 +2 +2 x=2

OpenStudy (anonymous):

x=2x/pi x=pi/4 x=7pi/4 x=pi/6 x=7pi/6

OpenStudy (loser66):

csc x = 2 \(\dfrac{1}{sinx} =2\) now switch, sinx = 1/2 x = pi/6

OpenStudy (anonymous):

ok thanks so much they should help me with some more of my problems :)